Undergraduate Physics Thesis Proposal

Matthew Gray

December 19, 1996

Introduction

A system of intrabody signalling has been developed in the Physics and Media group at the Media Lab. This system, referred to as “Personal Area Networking” (PAN) was developed in prototype by Tom Zimmerman. The system works by electrostatically driving the body through electrodes which are capacitavely coupled to the body.

By modulating this signal, data can be effectively transmitted through the body. Currently, prototype versions of this system exist, but the understanding of the constraints of the system are limited. The system is currently limited to a few geometries and there is very little understanding of the noise sources and limitations.

Description of Proposed Work

The goals of this thesis are to develop a more reliable prototype system and develop a numerical simulation of the system. Given a more flexible and reliable prototype, a variety of data encoding schemes can be tried as well as collecting experimental data on a variety of geometries. Specific encoding strategies to be tried include traditional AM and FM schemes as well as some more sophisticated techniques, so far to specifically include differential phase shift keying. By developing a detailed numerical model of the system, we hope to better understand the limits of PAN, in terms of the noise features of the system and the effect of varied geometries.

While the development of the prototype system will be valuable, it will not be the primary focus of this thesis. This thesis will focus on the development of the mathematical model for simulation of PAN. The full model will incorporate a successive over-relaxation (SOR) electric field solver, a circuit solver, and thermal noise model for the human body.

This work will be completed for graduation in June, 1997 under the supervision of Dr. Neil Gershenfeld in the Physics and Media Group of the MIT Media Lab.